OpenStudy (anonymous):

I'm having trouble solving this: Find an equation of the line containing the given pair of points. The points are: (1/4, -1/2) and (3/4,3). If you could really explain the steps when multiplying the fractions I'd really appreciate it. Thanks :)

OpenStudy (ghazi):

equation of line passing though two points is given by \[(Y-Y _{1})= \frac{ (Y _{2}-Y _{1}) }{ (X _{2}-X _{1}) } *(X-X _{1})\]

OpenStudy (ghazi):

lets say X1=1/4 , Y1= -1/2 and X2= 3/4 and Y2=3 , now plug in the values you will get an equation after a bit of simplification :D

OpenStudy (anonymous):

Thank you :3 I was able to figure it out with that. I wasn't using -1/2

OpenStudy (ghazi):

oh no problem :)